OpenStudy (anonymous):

x+2y(is greater than or equal to)-9: i need points for the linar equation

OpenStudy (wolfe8):

Eventhough it's an inequality, you can treat it like an equation and make this in the straight line general form. You you understand that part?

OpenStudy (anonymous):

not really..... :/

OpenStudy (wolfe8):

Do you know the form y=mx+b ?

OpenStudy (anonymous):

yes i do, is that what i use to get the two points?

OpenStudy (wolfe8):

No but you can get the gradient and y-intercept which should be enough to make a straight line.

OpenStudy (anonymous):

which would be -7??

OpenStudy (anonymous):

im sorry im so terrible at math!

OpenStudy (wolfe8):

Alright here I'll show you.

OpenStudy (anonymous):

omg i love you

OpenStudy (wolfe8):

So you have \[x+2y \ge-9\]As I said you can treat it like an equation so can make it so that you get the general straight line form y=mx+b So you subtract both sides with x and divide everything by 2 and get:\[y \ge-\frac{ 9 }{ 2 }-\frac{ 1 }{ 2 }x\] which can be rewritten as: \[y \ge-\frac{ 1 }{ 2 }x-\frac{ 9 }{ 2 }\] And this is similar to the form y=mx+b where m is the gradient and b is the y-intercept :) Can you make the straight line from here?

OpenStudy (anonymous):

ahsdlkjfeoij i hate math!!! but thank you!! um... does y=-9?

OpenStudy (anonymous):

idk! we just learned this today, and i was completely zoned out during class

OpenStudy (wolfe8):

The y-intercept? No it is -9/2 from the last equation I wrote :) It's ok you can learn. So from the last equation, you can see that the gradient, m, is -1/2 and the y intercept -9/2

OpenStudy (anonymous):

thank you so much! what ever your name is!! one last queestion.. if it is just greater than, is it a straight line? or a dotted line?

OpenStudy (wolfe8):

Ah so if has 'or equal to' it is a full line. If not, it's a dotted line.

OpenStudy (anonymous):

i love you, you are my new best friend

OpenStudy (wolfe8):

Haha you're welcome. My name is Wolfe btw :)

OpenStudy (anonymous):

im Hannah! obviously:) thanks again!